b. What was the increase in retained earnings for the year? Amigo Software Inc. has total assets of $889,000, current liabilities of $192,000, and long-term liabilities of $154,000. There is $87,000 in preferred stock outstanding. Thirty thousand shares of common stock have been issued. 4. a. Compute book value (net worth) per share. b. If there is $56,300 in earnings available to common stockholders and the firms stock has a P/E of 23 times earnings per share, what is the current price of the stock? c. What is the ratio of market value per share to book value per share? (Round to two places to the right of the decimal point.)
Answer #1

Question 4

a. For this we use the accounting equation: Assets = liabilities +equity

889,000 = 192,000+ 154,000 + 87,000 + Common Stock

Common Stock = 456,000

Book Value per share = 456,000/Number of shares 456,000/30,000 = $15.20

Book value per share =$15.20

b. Earnings per share = 56,300/30,000 = $1.8767/share

P/E = 23

P = 23* E = 23*1.8767

P = $43.1633

Current price of the stock = $43.16

c. Market Value to Book Value = 43.1633/15.20 = 2.84

Market Value to Book Value = 2.84